OpenStudy (anonymous):

solve the equation below for x. 100=25e^3x

OpenStudy (brimblings):

100/25 = e^3x 4=e^3x this implies that ln(4)=3x then solve to find x

OpenStudy (anonymous):

A.=ln4/3 B.x=ln100/3*ln25 C.x=4/x^3 D.x=ln4-3

OpenStudy (brimblings):

A

OpenStudy (anonymous):

thank you so much :)